OpenStudy (anonymous):

can someone help with 2nd order ODE homogeneous equation..

OpenStudy (anonymous):

2x"-7x'+6x=0 x'=dx/dt x"=d^2/dt^2

OpenStudy (amoodarya):

char equation is 2r^2-7r+6=0 so r=1.5 ,2 now x = a exp(1.5 t ) +b exp(2 t)

OpenStudy (anonymous):

i got to that part...and I get a and b equal zero...where am I going wrong?

OpenStudy (anonymous):

initial conditions...

OpenStudy (anonymous):

x(0)=0 x'(0)=0

OpenStudy (amoodarya):

apply the initial condition

OpenStudy (amoodarya):

x = a exp(1.5 t ) +b exp(2 t) x'=1.5 a exp(1.5 t) +2b exp(2 t) a,b goes to zero are you sure about initial conditions?

OpenStudy (anonymous):

yes and looking back, I just remembered the professor asking us what is so particular about them both equaling zero...what does it mean?

OpenStudy (anonymous):

nevermind, classmate has these IC: x(0)=1 x'(0)=0 but now I get -B+B=1...doesn't make sense

OpenStudy (amoodarya):

what grade r u in ? is it for ms degree ode ?

OpenStudy (anonymous):

senior engineer undergrad

OpenStudy (usukidoll):

what's the question?

OpenStudy (amoodarya):

x(0)=1 x'(0)=0 are those initial condition ?

OpenStudy (usukidoll):

^ whatever he said

OpenStudy (anonymous):

@amoodarya yes

OpenStudy (usukidoll):

because after you find your x = a exp(1.5 t ) +b exp(2 t) you have to apply x(0)= 1 first then take the derivative of x = a exp(1.5 t ) +b exp(2 t) and apply x'(0) =0

OpenStudy (amoodarya):

x = a exp(1.5 t ) +b exp(2 t) x(0)=1 --> a+b=1 x'=1.5 a exp(1.5 t) +2b exp(2 t) x'(0)=0 --> 1.5a +2b=0 now find a,b

OpenStudy (usukidoll):

I normally use c1 and c2 for these type of problems... easier to see x = c1e^1.5t + c2e^2t

OpenStudy (anonymous):

thank you, i'm getting all my writings mixed up

OpenStudy (usukidoll):

so I assume that I'm dealing with a x(t) = c1e^1.5t + c2e^2t x(0) = 1 1 = c1+ c2

OpenStudy (usukidoll):

e^ 0 is 1 ^^ but multiplied by c1 and it's c1

OpenStudy (usukidoll):

so we need to product rule for x(t)

OpenStudy (amoodarya):

a+b=1 1.5a+2b=0 if u solve a=4 , b=-3

OpenStudy (usukidoll):

let me double check that x(t) = c1e^1.5t + c2e^2t x'(t) = c1 1.5e^1.5t + c2 2e^2t

OpenStudy (usukidoll):

x'(0) = 0 0 = 1.5 c1 +2c2

OpenStudy (usukidoll):

1 = c1+ c2 0 = 1.5 c1 +2c2

OpenStudy (usukidoll):

subsitution method... or matrix method.. would be easier.

OpenStudy (anonymous):

@amoodarya can you help with the non homogeneous part now?

OpenStudy (amoodarya):

yes

OpenStudy (anonymous):

ok I will post in a bit and tag you..

OpenStudy (usukidoll):

uh oh... use subsitution method to solve for c1 and c2.

OpenStudy (usukidoll):

4 and - 3 is right

OpenStudy (anonymous):

ok so the whole problem was.. 2x"-7x'+6x=sin2t i solved for the homogeneous part, now i need to solve for the particular solution..the professor showed us how to do it if the right side was a constant but you can see that it is not so I'm not sure what I'm suppose to do now @amoodarya

OpenStudy (usukidoll):

oh these.. undetermined coefficents.

OpenStudy (usukidoll):

yes you do have to find the Yh first which is what we did... we just need Yp

OpenStudy (anonymous):

thank you for your input also. it sounds familiar..

OpenStudy (usukidoll):

there's a format for the right hand side... there's no e^xsinx or e^x cosx so it may make life easier... sinwx coswx has the format of Asinwx + Bcoswx where w = the number

OpenStudy (usukidoll):

sin2t = Asin2t + Bcos2t

OpenStudy (anonymous):

but they are new constants?

OpenStudy (usukidoll):

Yp = Asin2t + Bcos2t find the first and second derivatives of this equation... then plug it in to 2x"-7x'+6x=sin2t now if you have a 0 = sin 2t that means that our initial Yp fails and we need to add a mulitplier

OpenStudy (usukidoll):

so find the derivatives of Yp and plug it in to the equation

OpenStudy (usukidoll):

then collect all of your A's and B's.

OpenStudy (anonymous):

you mean plug it in x" and x' after I find the derivatives?

OpenStudy (usukidoll):

:) yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so far I have -74sin2t-32cos2t+24e^(3/2t)-18e^(2t)=sin2t

OpenStudy (anonymous):

gotta go for a bit brb

OpenStudy (usukidoll):

ummm whoa no no no no no no no no no that's not what I meant X*(

OpenStudy (usukidoll):

you have to plug in the derivatives and the original Yp in 2x"-7x'+6x=sin2t and I recommend that you use y instead 2y''-7y'+6y = sin2t

OpenStudy (anonymous):

o oops

OpenStudy (anonymous):

I'm following...but i forgot to ask why sin2t=Asin2t+Bcos2t? just a rule i need to remember?

OpenStudy (usukidoll):

yeah it's a format.... you need to remember for sinwt and coswt it's Asinwt+Bcoswt where w = the number your assigned if applicable.

OpenStudy (usukidoll):

there's four of them... e^cx c = exponent number x^n n= degree number the one I typed earlier and another one with the one I typed earlier with e^cx